OpenStudy (anonymous):

Three cards are drawn without replacement from a well-shuffled deck of 52 playing cards. What is the probability that the third card drawn is a heart?

OpenStudy (kropot72):

If a heart is not drawn on the first and second draws, P(heart) = 13/50. If a heart is drawn on the first or second draws, P(heart) = 12/50. If hearts are drawn on the first and second draws, P(heart) = 11/50.